OpenStudy (anonymous):

1. Find the absolute maximum and minimum values of the function f(x)=12+4x-x^2 in the interval [0,5]

ganeshie8 (ganeshie8):

find the derivative, and, set it equal to 0

ganeshie8 (ganeshie8):

that gives u critical points where max/min values can occur

OpenStudy (anonymous):

yes, I did that. so it'll only be x=2?

ganeshie8 (ganeshie8):

x = 2 is one possibility where max/min value can occur

ganeshie8 (ganeshie8):

other possibilities are x = 0 and x = 5

OpenStudy (anonymous):

yes yes :) I'm not sure but I got 12, 7 and 16 :)

ganeshie8 (ganeshie8):

so, here is the thing : Absolute maximum = Maximum of {f(2), f(0), f(5)} Absolute minimum = Minimum of {f(2), f(0), f(5)}

OpenStudy (anonymous):

I thought there can be only a maximum of 4 values to choose from? :)

ganeshie8 (ganeshie8):

so : Absolute maximum = Maximum of {16, 12, 7} = 16 occurs when x = 2

ganeshie8 (ganeshie8):

Absolute manimum = Minimum of {16, 12, 7} = 7 occurs when x = 5

ganeshie8 (ganeshie8):

since you're given BOUNDARY, both max and min will be there.

ganeshie8 (ganeshie8):

if you're not given any boundary, then oly max will be there for the given function

OpenStudy (anonymous):

ohh :)

OpenStudy (anonymous):

can you please help me with 2 more? it's harder :(

ganeshie8 (ganeshie8):

so to find absolute max/mins : u need to check both 'local max/mins' and the 'boundary values'

ganeshie8 (ganeshie8):

sure :)

OpenStudy (anonymous):

thankyou so much 2. Find the absolute maximum and minimum values of the function \[g(x)=(x^{2}-1)^{3}\] in the interval [-1,2]

ganeshie8 (ganeshie8):

step1 : take the derivative, set it equal to 0, and solve x

OpenStudy (anonymous):

okay, Im going to try it :)

OpenStudy (anonymous):

chain rule? :O

ganeshie8 (ganeshie8):

yess

OpenStudy (anonymous):

I'm now on \[6x(x ^{2}-1)^{2}=0\] and I really don't know what to do with it :"(

ganeshie8 (ganeshie8):

solve x

ganeshie8 (ganeshie8):

use zero-product-property

OpenStudy (anonymous):

oh okay, I think I get it

OpenStudy (anonymous):

is it x=0,1 ?

ganeshie8 (ganeshie8):

u left one value :)

ganeshie8 (ganeshie8):

\(\large 6x(x^2-1) = 0 \) \(\large \implies x = 0 \) or \(\large \implies x^2-1 = 0 \)

ganeshie8 (ganeshie8):

\(\large x^2-1 = 0 \implies x^2 = 1 \implies x = \pm 1\)

ganeshie8 (ganeshie8):

so the local max/mins are : x = 0, 1, -1

ganeshie8 (ganeshie8):

okay ?

OpenStudy (anonymous):

oh okay :)

ganeshie8 (ganeshie8):

step2 : evaluate the given function at 1) local max/mins and 2) boundary

OpenStudy (anonymous):

I'm going to solve for it now :)

ganeshie8 (ganeshie8):

okie

OpenStudy (anonymous):

max:9 min:0?

ganeshie8 (ganeshie8):

basically u need to evaluate below : g(0) g(-1) g(1) g(2)

ganeshie8 (ganeshie8):

let me check..

ganeshie8 (ganeshie8):

\(\large g(x) = (x^2-1)^3\) g(0) = -1 g(-1) = 0 g(1) = 0 g(2) = 27

ganeshie8 (ganeshie8):

so, minimum value = -1, occurs when x = 0

OpenStudy (anonymous):

oh no. haha

ganeshie8 (ganeshie8):

maximum value = 27, occurs when x = 2

OpenStudy (anonymous):

sorry. my bad. I wrote here (3)^3, thought it was 3 x 3

ganeshie8 (ganeshie8):

haha i guessed that :)

ganeshie8 (ganeshie8):

graph looks like that, it may be exciting to see that the graph has minimum value at x = 0 ... :)

OpenStudy (anonymous):

oh cool graph. haha sorry, another careless mistake. I wrote -1 here and I encircled it but I didn't saw it :(

OpenStudy (anonymous):

last item number :) I find this really difficult because it involves trigonometry :(

ganeshie8 (ganeshie8):

oh lets see -.-

OpenStudy (anonymous):

3. Find the absolute maximum and minimum values of the function h(x)=2cost + sin2t in the interval\[[0,\pi/2]\]

ganeshie8 (ganeshie8):

step1 : take the derivative, set it equal to 0, and solve x

ganeshie8 (ganeshie8):

*or t, or watever the variable is

OpenStudy (anonymous):

uhm, how do you find the derivative when its sin2t again? haha I completely forgot about it sorry

ganeshie8 (ganeshie8):

chain rule to the rescue !

OpenStudy (anonymous):

ohh haha

ganeshie8 (ganeshie8):

\(\large \frac{d}{dt}[\sin(2t)] = \cos(2t) * \frac{d}{dt}(2t) = \cos(2t) * 2 = 2\cos(2t)\)

ganeshie8 (ganeshie8):

first, find the derivative of h(t) and set it equal to 0 later we can see how to solve it :)

OpenStudy (anonymous):

okay :) i'll try my best

OpenStudy (anonymous):

but ...

ganeshie8 (ganeshie8):

good, give it a try :)

ganeshie8 (ganeshie8):

\(\large h(t) = 2\cos(t) + \sin(2t)\) \(\large h'(t) = -2\sin(t) + 2\cos(2t)\)

ganeshie8 (ganeshie8):

set it equal to 0

ganeshie8 (ganeshie8):

\(\large h'(t) = 0\) \(\large -2\sin(t) + 2\cos(2t) = 0 \) \(\large -\sin(t) + \cos(2t) = 0 \) \(\large -\sin(t) + 1-2\sin^2t= 0 \) \(\large 2\sin^2t + \sin t -1 = 0 \)

OpenStudy (anonymous):

but how? it has sin and cos? :"(

ganeshie8 (ganeshie8):

we got rid of cos's already :) go thru it and see if it looks okay... let me knw if something doesnt make sense ..

ganeshie8 (ganeshie8):

we still need to solve to \(t\) ...

OpenStudy (anonymous):

is 1 -2sin2t an identity I should know about? :O

ganeshie8 (ganeshie8):

good question :) yes u need to memorize below trig identities : \(\large \cos 2t = 1-2\sin^2 t\) \(\large \cos 2t = 2\cos^2 t-1\)

OpenStudy (anonymous):

the 2nd one, I think that should be sine? :)

ganeshie8 (ganeshie8):

nope. its cos oly..

OpenStudy (anonymous):

really? are there any more identities? :)

ganeshie8 (ganeshie8):

one more important one is : \(\large \sin^2 t + \cos^2t = 1\)

ganeshie8 (ganeshie8):

another important one is : \(\large \sin 2t = 2 \sin t \cos t\)

ganeshie8 (ganeshie8):

those four identities will do the job most of the time...

OpenStudy (anonymous):

oh okaay thankyou so much and now back to solving :)

ganeshie8 (ganeshie8):

yeah where are we..

ganeshie8 (ganeshie8):

\(\large 2\sin^2t + \sin t -1 = 0 \) we are at solving this equation for \(t\)

OpenStudy (anonymous):

huh? not sin(t)? O_O

ganeshie8 (ganeshie8):

nope, we want to find \(t\) that gives the min/max values

OpenStudy (anonymous):

ohh

ganeshie8 (ganeshie8):

... for the given function h(t)

ganeshie8 (ganeshie8):

\(\large 2\sin^2t + \sin t -1 = 0 \) if u replace \(\sin t\) wid \(x\), clearly its a quadratic equation in x's right ?

OpenStudy (anonymous):

yeah :)

ganeshie8 (ganeshie8):

so, can u factor it ha ? :)

OpenStudy (anonymous):

(2x+1)(x-1)

OpenStudy (anonymous):

\[(2sint +1 ) (sint -1)\]

OpenStudy (anonymous):

is it correct? :)

ganeshie8 (ganeshie8):

partially correct... ^careful about the signs

ganeshie8 (ganeshie8):

\(\large 2\sin^2t + \sin t -1 = 0 \) \(\large 2\sin^2t + 2\sin t - \sin t -1 = 0 \) \(\large 2\sin t(\sin t + 1) - 1( \sin t + 1) = 0 \) \(\large 2\sin t(\sin t + 1) - 1( \sin t + 1) = 0 \) \(\large (2\sin t - 1)( \sin t + 1) = 0 \)

ganeshie8 (ganeshie8):

okay ?

OpenStudy (anonymous):

sorry :)

ganeshie8 (ganeshie8):

np :) solve t

OpenStudy (anonymous):

\[\sin(t)=\frac{ 1 }{ 2 },-1\]

ganeshie8 (ganeshie8):

perfect ! solve t

ganeshie8 (ganeshie8):

\(\large (2\sin t - 1)( \sin t + 1) = 0 \) \(\large \implies 2\sin t - 1 = 0\) or \(\large \sin t + 1 = 0\) \(\large \implies \sin t = \frac{1}{2}\) or \(\large \sin t = -1\) \(\large \implies t = \frac{\pi}{3}\)

ganeshie8 (ganeshie8):

so oly one value we got for \(t\) after so much work -.-

ganeshie8 (ganeshie8):

btw, we had to discard -1 cuz that gives t = pi, but the question is asking us to find in the boundary (0, pi/2)